OpenStudy (gokuporter):

Enter the value for that makes the given equation true. 3x−5=−8(6+5x)

pooja195 (pooja195):

3x−5=−8(6+5x) Start by distributing 3x−5=(−8)(6)+(−8)(5x) What equation do we have left when we do this?

OpenStudy (gokuporter):